2. What fraction of oxygen bound in the lungs is released in peripheral tissues for each hemoglobin? Which hemoglobin is better at transporting oxygen? Assume that the lung pO2 = 100 torr and that the peripheral tissue pO2 is 20 torr. Based on the shape of the saturation curve shown above, describe the effect of the Asp99Asn substitution on the cooperativity of oxygen binding? Is the cooperatively positively or negatively impacted by this mutation? 3. 4. In normal deoxyhemoglobin (hemoglobin with no oxygen bound, the T state) the β- subunit Asp 99 forms a hydrogen bond with the phenolic hydroxyl of Tyr 42 from the a-subunit. This hydrogen bond stabilizes the subunit interface of deoxyhemoglobin. Has the Asp99Asn mutation made the deoxy form more, or less, stable? Provide a plausible (structural) explanation. Based on your answer to Q4, provide a general mechanism by which mutations in hemoglobin could effects it s ability to transport oxygen. 5.
media%2F531%2F5314a885-cf11-42be-8fb6-b2
Answer #1

1. The hemoglobin Kempsey has higher affinity for oxygen. The P50 for oxygen at pH 7.2 is 10% of haemoglobin A.

P50 of haemoglobin is the oxygen tension at which 50% of haemoglobin is saturated. In the oxyhaemoglobin dissociation curve depicted, Hb A has a P50 value of 30 Torr, while that of Hb Kempsey is 10 Torr. Hence, Hb Kempsey is saturated at a much lower oxygen tension than Hb A. Hb Kempsey has a much higher affinity for oxygen.

2. The partial pressure of lungs is 100 Torr and of peripheral tissue. The difference in fraction saturated between the two components will give the fraction of oxygen bound to the lungs that is release in peripheral tissue. Consider this to be X.

X for HbA:

Lung fraction saturated = 0.98 (Lungs have 98% oxygen saturation)

Peripheral Tissue= 0.25 (Peripheral tissue have 25% oxygen saturation)

Both values are also seen in graph.

X for HbA= 0.98-0.25= 0.73

This is the amount of oxygen present in systemic veins.

X for Hb Kempsey:

Lung fraction saturated = 0.98

Peripheral Tissue= 0.85

X for Hb Kempsey= 0.98-0.85= 0.13

Hence, less oxygen is released by haemoglobin Kempsey into the tissues than Hb A.

3. Haemoglobin is a tetramer of two alpha and two beta subunits liked by a heme molecule. Binding of one molecule of oxygen to the monomeric subunit of Hb, causes other oxygen molecules to bind to other three sites in Hb. This feature is known as co-operativity of oxygen binding. Such co-operativity is due to a structural or conformational change in Hb structure. This creates a hyperbolic curve for Hb. Hence, HbA has a sigmoidal oxygen dissociation curve. However, Hb Kempsey binds oxygen faster. Its dissociation curve is hyperbolic rather than sigmoidal. Hence, the co-operativity is positively affected by the mutation. Few oxygen molecules are bound to this Hb. Hence, dissociation will be faster.

4. Less oxygen is released by Hb Kempsey than HbA due to perturbation in the alpha 1 beta 2 interactions that prevents formation of hydrogen bonds between β99 Asp and α42 Tyr. There are two states of Hb- R and T states. The T state has low affinity for oxygen while the R state has high affinity for oxygen. For a concerted mode of op-operativity, the Hb should be either in T or R state. Hb shifts rapidly between these two states to transport oxygen. The bonds between β99 Asp and α42 Tyr normally stabilize deoxyHb low oxygen affinity T state. There is a shift of quaternary equilibrium to oxygenated R state of Hb due to no bond formation between the Asp and Tyr residues. Hence, the T form, deoxyHb, is destabilized. As a result, less oxygen is released to peripheral tissues.

As per Chegg’s rule, only one question is mandatorily answered.